# ToolHaystack: Stress-Testing Tool-Augmented Language Models in Realistic Long-Term Interactions

### v1

@@ Abstract
Beong-woo Kwak1  Minju Kim1  Dongha Lim1  Hyungjoo Chae1 Dongjin Kang1Sunghwan Kim1Dongil Yang1Jinyoung Yeo1 1Yonsei University {bwoo.kwak,jinyeo}@yonsei.ac.kr
@@ 1 Introduction
Recent breakthroughs in Large Language Models (LLMs) have transformed their role toward LLM agents; they can execute real-world tasks such as managing financial transactions and scheduling appointments (Yao et al., 2022; Shinn et al., 2023). Tool-augmented language model (TALM) benchmarks (Qin et al., 2023; Li et al., 2023; Chen et al., 2024) have played a critical role in evaluating tool-use capabilities based on simplified instruction-following setups where they autonomously execute external tools to address user goals.
@@
Despite these advancements, the scope of most existing work is still limited short interaction, which fail to capture the behavior of LLM agents in long-term interactions. The queries of most existing benchmarks are simple (e.g.“Book a restaurant named ….”) which can be resolved within one or two exchanges. In contrast, many real-world tasks, such as travel planning or project coordination, cannot be resolved in short interactions (Wu et al., 2025; Maharana et al., 2024) or steps (Hayati et al., 2025; Press et al., 2023). For example, the user may want to perform tasks based on previous requests without restating all the details (e.g.“Cancel the reservation you made for my birthday.”) or request to collaborate with the agents across long time frame (e.g.“Help me writing a research paper.”). These tasks inherently require interactions that unfold and might evolve over time.
@@
While recent efforts have extended test scenarios through multi-turn interactions (Li et al., 2023) and API response simulations (Yan et al., 2024; Zhong et al., 2025), these approaches fall short in capturing the full complexity of long-term interactions. In real-world contexts, continuous exchanges between the user, the agent, and the external environment introduce various contextual noises—such as interleaved task flows (Castillo-Bolado et al., 2024), redundant information (Bai et al., 2024), and shifting user goals (Yu et al., 2025)—that accumulate in the agent’s context (Fig.1). The current lack of benchmarks that rigorously evaluate TALMs’ robustness under such conditions limits their applicability to real-world scenarios, especially for long-term use cases like personal assistants.
@@
In this paper, we aim to evaluate TALMs in realistic long-term interaction scenarios. To this end, we introduce ToolHaystack, a realistic and composable test suite for TALMs. Specifically, we define long-term agent context as a highly noisy context where multiple user goals and partial task execution histories are entangled in a single context. Test instances are structured to simulate natural task flows by interleaving task-relevant ("needles") and irrelevant contexts ("haystack") [^2] We draw the analogy from the classic needle in a haystack problem (Kamradt, 2023) to emphasize the noisy nature of long-term interaction., posing unique challenges of identifying key information among distractions accumulated over time. Based on the structure, ToolHaystack simulates six realistic scenarios derived from three core challenges (context recall, information shift, and missing context), each with two difficulty levels.
@@
We systematically evaluate 14 highly capable LLMs with tool-use capabilities, covering both open-source and proprietary models on ToolHaystack. Our findings reveal that even highly capable LLMs with advanced long-context modeling (Bai et al., 2024; An et al., 2024) still struggle to maintain consistency and robustness under long-term scenarios—suggesting that contemporary benchmarks may overestimate their readiness and reliability for real-world deployment. Beyond general performance, we also conduct detailed ablation studies to uncover what factors influence long-term success and where current models fail.
@@
  • We are the first to explicitly formalize and highlight the critical challenge of long-term interaction robustness for TALMs.
@@ 2 Related Work
To better approximate real-world conditions, benchmarks such as ToolDial (Shim et al., 2025), HammerBench (Wang et al., 2024), and $\tau$-Bench (Yao et al., 2025) introduce more naturalistic, multi-turn dialogues. Recent efforts like RoTBench (Ye et al., 2023) and FAIL-TALM (Treviño et al., 2025) examine robustness under adversarial or erroneous conditions. However, these benchmarks remain largely constrained to single-session or short-horizon tasks, offering limited insight into TALM behavior over extended interactions.
@@
Recent benchmarks such as MMTB (Yu et al., 2025), BFCL (Yan et al., 2024), and ComplexFuncBench (Zhong et al., 2025) have extended TALM evaluation into longer contexts and compositional tool use. However, they often rely on clean task structures with highly coherent subtasks (< 5) or simulate long context by concatenating tool outputs.
In contrast, ToolHaystack provides a long-context TALM benchmark with interleaved task flows and contextual noise. This exposes novel failure modes that are underexplored in prior evaluations. Tab.1 compares ToolHaystack with existing TALM benchmarks. Further discussions on related work are presented in Appendix A.
@@ 3 Rethinking Evaluation of TALMs
Most LLM benchmarks are rapidly saturated by frontier models (Sirdeshmukh et al., 2025; Phan et al., 2025). Likewise, previous TALM evaluation frameworks might have limited lifespans. This underscores a constant need for more realistic, challenging and adaptive evaluation frameworks that can keep pace with model advancements. We highlight that long-term interactions go beyond simple extension of standard multi-turn evaluation.
@@ 4.1 Test Structure
To address this issue, we design a novel test structure for ToolHaystack, inspired by Needle-In-A-Haystack (NIAH) dataset (Kamradt, 2023). The key idea is to interleave different task sessions as a context instead of isolating task sessions (Fig.1). In ToolHaystack, needles are defined as parts of the context that are essential to address the target task. The term haystack refers to the remaining (usually irrelevant) text, that acts as a distractor. This modularized structure enables us to control the difficulty of test instances by manipulating how we compose this haystack. For example, increasing the number of task sessions in the haystack can naturally increase the task difficulty.
@@ 4.2 Test Scenarios
To rigorously evaluate the capabilities of TALMs in long-term scenario, we construct a taxonomy of scenarios based on three essential abilities reflecting realistic long-term tool-use challenges: Context Recall, Information Shift, and Missing Context. Each scenario is further divided into two levels of difficulties (i.e.simple and complex), yielding 6 scenarios in total.
@@ 5 Building ToolHaystack
Figure 3: Overview of our three-stage dataset generation pipeline.
@@ 5.1 Tool Set Collection
We begin by collecting over 13K real-world APIs from RapidAPI (RapidAPI, 2025), spanning 42 categories. Each API comes with a description of functionality and argument schema. We first sample a target API to be invoked in the target session from the curated toolset.
@@ 5.2 Scenario-Grounded Tool Sequence Construction
Each argument’s dependency is identified using semantic matching (SentenceBERT (Reimers and Gurevych, 2019)) and fine-grained LLM verification to ensure the semantic relevance and compatibility. We annotate each tool call with realistic argument values using LLM prompting. Specifically, a tuple of 4 elements is given to the model i.e., target API, needle APIs, a scenario type (e.g., CR, IS, and MC), situational context. Here, situation context is introduced to enhance the authenticity and consistency of the argument values. We pre-constructed a set of situational contexts for each goal APIs based on the characteristics and requirements of each target API and randomly use one of them as a situational context.
@@ 6.1 Experimental Setup
Call $\text{Call Acc}=\frac{\sum_{i=1}^{N}c_{i}}{\sum_{i=1}^{N}n_{i}}$ (1)
@@
To provide deeper insight, we generate a multi-turn counterpart of our dataset generated by standard self-instruct approach previously used in the literature (Qin et al., 2023). We use the same goal APIs for generating short multi-turn tool-use instances. All necessary information to complete the task is revealed within a few turns, without any distractions.
@@ 6.2 Main Results
Closed-source models consistently outperform open-source models under noisy settings. While general-purpose open-source models (Qwen2.5, DeepSeek-V3) demonstrate more stable performance across scenarios, they still lag behind top proprietary models. Notably, open-source finetuned models like ToolACE and watt-tool achieve generally achieve poor accruracy of less than 10%, despite their high BFCL scores. We further examine this by extending the comparison between multi-turn and long-term tool-use performance using BFCL benchmark in the following section.
@@ 6.5 Error Analysis through the Lens of Hallucination
Closed and General-purpose models exhibit highly skewed error distributions, with over 75% of their failures attributable to Out-of-context hallucinations in MC-Easy and similarly high proportions in MC-Hard (63% and 77%). We find that several models tend to copy default values of missing parameter or filling out randomly generated strings without proper verification of contextual sufficiency. Finetuned TALMs, in contrast, show a different error distribution. In MC-Easy, TALMs reduce Out-of-context hallucinations to 43.39% and In-context hallucinations remain low (7.94%), while API errors account for nearly half (48.68%) of the total errors, suggesting that TALMs are more cautious about fabricating arguments outright from missing context (lower Out-of-context) but often misidentify the appropriate tool (API error), particularly under harder conditions. See Appendix D.1 for detailed case studies illustrating model behaviors.
@@ A.1 Tool-Augmented Language Models
Large Language Models (LLMs) augmented with external tools have emerged as a promising paradigm for overcoming inherent limitations in static parametric memory (Komeili et al., 2022) and enabling real-time interaction with the external world (Yang et al., 2024; Chae et al., 2025). The term tool encompasses a broad range of functionalities, from simple calculators and Python interpreters (Cai et al., 2024) to complex API and function calls (Liu et al., 2024). In this work, we focus specifically on the latter—LLMs interfacing with external APIs or invoking function calls to perform tasks.
@@
Early efforts concentrated on enabling LLMs to call tools via prompting and reasoning frameworks. For instance, Parisi et al. (2022) introduce the concept of Tool Augmented Language Models (TALM), demonstrating that an LLM can iteratively interact with non-differentiable tools through textual interfaces. Similarly, the ReAct framework (Yao et al., 2022) combined chain-of-thought reasoning with action directives, allowing an LLM to plan steps and invoke tools (e.g., web search or calculators) in a conversational loop. Subsequent research has shown that fine-tuning LLMs on tool-use data can significantly enhance their efficacy. Toolformer (Schick et al., 2023), for example, is a fine-tuned model with self-generated tool API calls inserted into its training data, enabling it to decide when and how to use calculators, search engines, translators, and other APIs.
@@
With the rising demand for practical task-solving capabilities, recent proprietary LLMs have evolved to support tool or function calling as a core interface. Starting with GPT-4’s function calling API (OpenAI, 2023), OpenAI has progressively enhanced tool interoperability in their models, culminating in GPT-4.1 and GPT-4o, which natively support multimodal input, extended context windows, and robust tool invocation within conversation flows. Similarly, models like Claude 3.5 (Anthropic, 2024) and Gemini 1.5 (Team et al., 2024) offer built-in support for external API access and tool orchestration, enabling seamless integration of search, calculator, code execution, and retrieval systems in their agent pipelines. These developments mark a paradigm shift from pure text generation to real-world interaction, where tool use becomes a first-class citizen in the model’s interface. Alongside model-level enhancements, ecosystem-level tool agents have also proliferated. For example, OpenAI’s ChatGPT integrates plugin systems for third-party APIs and file browsing, while Google’s Bard/Gemini and Anthropic’s Claude provide tool-use scaffolding through memory and workflow APIs. These systems often combine planning, execution, and memory modules, highlighting the emerging need for robust, API-centric LLM benchmarks and evaluation methods.
@@
Recognizing the performance gap between open-source LLMs and proprietary models in tool calling, recent efforts have focused on equipping smaller or open LLMs with competitive tool-use abilities. For instance, Gorilla (Patil et al., 2023) fine-tunes LLaMA models on a large corpus of API documentation and call traces, leveraging a retriever-generator architecture to improve accuracy and reduce hallucination. ToolLLM (Qin et al., 2023) adopts a reasoning-guided planning strategy that explores multiple API invocation paths via a depth-first search over possible tool sequences. Salesforce’s xLAM (Zhang et al., 2024) provides a family of open LLMs specifically tuned for tool usage, achieving strong performance on multi-tool benchmarks. Other open initiatives such as ToolAlpaca (Tang et al., 2023), AnyTool (Du et al., 2024), and ToolBridge (Jin et al., 2024) introduce simulated training pipelines or meta-learning objectives for improving tool generalization in smaller models. These efforts aim to bridge the gap in real-world utility and provide a competitive open-source alternative to commercial tool-augmented LLM systems.

### v2

@@ Abstract
Beong-woo Kwak1   Minju Kim1   Dongha Lim2   Hyungjoo Chae1 Dongjin Kang1Sunghwan Kim1Dongil Yang1Jinyoung Yeo1 1Department of Artificial Intelligence, Yonsei University 2Department of Computer Science & Engineering, Yonsei University {bwoo.kwak,jinyeo}@yonsei.ac.kr
@@ 1 Introduction
Recent breakthroughs in Large Language Models (LLMs) have transformed their role toward LLM agents; they can execute real-world tasks such as managing financial transactions and scheduling appointments (yao2022react; shinn2023reflexion). Tool-augmented language model (TALM) benchmarks (qin2023toolllm; li2023apibank; chen2024t) have played a critical role in evaluating tool-use capabilities based on simplified instruction-following setups where they autonomously execute external tools to address user goals.
@@
Despite these advancements, the scope of most existing work is still limited short interaction, which fail to capture the behavior of LLM agents in long-term interactions. The queries of most existing benchmarks are simple (e.g.“Book a restaurant named ….”) which can be resolved within one or two exchanges. In contrast, many real-world tasks, such as travel planning or project coordination, cannot be resolved in short interactions (wulongmemeval; maharana2024locomo) or steps (hayati2025chain; press2023measuring). For example, the user may want to perform tasks based on previous requests without restating all the details (e.g.“Cancel the reservation you made for my birthday.”) or request to collaborate with the agents across long time frame (e.g.“Help me writing a research paper.”). These tasks inherently require interactions that unfold and might evolve over time.
@@
While recent efforts have extended test scenarios through multi-turn interactions (li2023apibank) and API response simulations (berkeley-function-calling-leaderboard; zhong2025complexfuncbench), these approaches fall short in capturing the full complexity of long-term interactions. In real-world contexts, continuous exchanges between the user, the agent, and the external environment introduce various contextual noises—such as interleaved task flows (castillo2024beyond), redundant information (bai2024longbench), and shifting user goals (yu2025multi)—that accumulate in the agent’s context (Fig.1). The current lack of benchmarks that rigorously evaluate TALMs’ robustness under such conditions limits their applicability to real-world scenarios, especially for long-term use cases like personal assistants.
@@
In this paper, we aim to evaluate TALMs in realistic long-term interaction scenarios. To this end, we introduce ToolHaystack, a realistic and composable test suite for TALMs. Specifically, we define long-term agent context as a highly noisy context where multiple user goals and partial task execution histories are entangled in a single context. Test instances are structured to simulate natural task flows by interleaving task-relevant ("needles") and irrelevant contexts ("haystack") [^2] We draw the analogy from the classic needle in a haystack problem (kamradt2023needle) to emphasize the noisy nature of long-term interaction., posing unique challenges of identifying key information among distractions accumulated over time. Based on the structure, ToolHaystack simulates six realistic scenarios derived from three core challenges (context recall, information shift, and missing context), each with two difficulty levels.
@@
We systematically evaluate 14 highly capable LLMs with tool-use capabilities, covering both open-source and proprietary models on ToolHaystack. Our findings reveal that even highly capable LLMs with advanced long-context modeling (bai2024longbench; an2024eval) still struggle to maintain consistency and robustness under long-term scenarios—suggesting that contemporary benchmarks may overestimate their readiness and reliability for real-world deployment. Beyond general performance, we also conduct detailed ablation studies to uncover what factors influence long-term success and where current models fail.
@@
  • We are the one of the pioneering works to explicitly formalize and highlight the critical challenge of long-term interaction robustness for TALMs.
@@ 2 Related Work
To better approximate real-world conditions, benchmarks such as ToolDial (shimtooldial), HammerBench (wang2024hammerbench), and $\tau$-Bench (yao2025taubench) introduce more naturalistic, multi-turn dialogues. Recent efforts like RoTBench (ye2023rotbench) and FAIL-TALM (trevino2025benchmarking) examine robustness under adversarial or erroneous conditions. However, these benchmarks remain largely constrained to single-session or short-horizon tasks, offering limited insight into TALM behavior over extended interactions. Some recent benchmarks such as MultiChallenge (sirdeshmukh2025multichallenge) have attempted to evaluate language models in multi-turn scenarios systematically. Compared to MultiChallenge, ToolHaystack differs significantly in problem space, failure types, and evaluation design. MultiChallenge focuses on instruction-following in relatively short and clean conversations ( 5 turns), with limited extent and control over noise structure. In contrast, ToolHaystack targets tool-use scenarios with an average of over 32 turns per instance, explicit argument chaining over long-horizon, and adjustable, scenario-specific noise injection. These settings better reflect real-world task complexity and allow us to uncover nuanced model brittleness under long-term, noisy interactions.
@@
Recent benchmarks such as MMTB (yu2025multi), BFCL (berkeley-function-calling-leaderboard), and ComplexFuncBench (zhong2025complexfuncbench), ACEBench (chen2025acebench), APIGen-MT (prabhakar2025apigen) have extended TALM evaluation into longer contexts and compositional tool use. However, they often rely on clean task structures with highly coherent subtasks (< 5) or simulate long context by concatenating tool outputs. In contrast, ToolHaystack provides a long-context TALM benchmark with interleaved task flows and contextual noise. This exposes novel failure modes that are underexplored in prior evaluations. Tab.1 compares ToolHaystack with existing TALM benchmarks. Further discussions on related work are presented in Appendix A.
@@ 3 Rethinking Evaluation of TALMs
Most LLM benchmarks are rapidly saturated by frontier models (sirdeshmukh2025multichallenge; phan2025humanity). Likewise, previous TALM evaluation frameworks might have limited lifespans. This underscores a constant need for more realistic, challenging and adaptive evaluation frameworks that can keep pace with model advancements. We highlight that long-term interactions go beyond simple extension of standard multi-turn evaluation.
@@ 4.1 Test Structure
To address this issue, we design a novel test structure for ToolHaystack, inspired by Needle-In-A-Haystack (NIAH) dataset (kamradt2023needle). The key idea is to interleave different task sessions as a context instead of isolating task sessions (Fig.1). In ToolHaystack, needles are defined as parts of the context that are essential to address the target task. The term haystack refers to the remaining (usually irrelevant) text, that acts as a distractor. This modularized structure enables us to control the difficulty of test instances by manipulating how we compose this haystack. For example, increasing the number of task sessions in the haystack can naturally increase the task difficulty.
@@ 4.2 Test Scenarios
To rigorously evaluate the capabilities of TALMs in long-term scenario, we construct a taxonomy of scenarios based on three essential abilities reflecting realistic long-term tool-use challenges: Context Recall, Information Shift, and Missing Context. Each scenario is further divided into two levels of difficulties (i.e., simple and complex), yielding 6 scenarios in total.
@@ 4.2.3 Missing Context (MC)
Figure 4: Overview of our three-stage dataset generation pipeline.
@@ 5.1 Tool Set Collection
We begin by collecting over 13K real-world APIs from RapidAPI (rapidapi), spanning 42 categories. Each API comes with a description of functionality and argument schema. We first sample a target API to be invoked in the target session from the curated toolset.
@@ 5.2 Scenario-Grounded Tool Sequence Construction
Each argument’s dependency is identified using semantic matching (SentenceBERT (reimers2019sentence)) and fine-grained LLM verification to ensure the semantic relevance and compatibility. We annotate each tool call with realistic argument values using LLM prompting. Specifically, a tuple of 4 elements is given to the model i.e., target API, needle APIs, a scenario type (e.g., CR, IS, and MC), situational context. Here, situation context is introduced to enhance the authenticity and consistency of the argument values. We pre-constructed a set of situational contexts for each goal APIs based on the characteristics and requirements of each target API and randomly use one of them as a situational context.
@@ 5.3 Dialogue Instantiation
5.4 Quality Assurance
We employ multiple frontier LLMs (GPT-4o, Claude 3.5, Gemini 2.5, LLaMA 3.3) during generation to reduce model-specific artifacts. Automated validation enforces constraints such as correct argument use, proper tool syntax, and scenario fidelity. A subset of generated dialogues undergoes a filtering process, where we prompt gpt-4o to strictly evaluate if the dialogue violates the rules for the target category. Moreover, to ensure the quality of ToolHaystack, we employed GPT-4o as a judge to automatically filter out dialogues that do not satisfy our scenario definitions. To validate the effectiveness of this automated filtering process, we sampled 60 dialogues (10 from each category) from the unfiltered dataset and compared the results against human annotations. The evaluation yielded a Matthews Correlation Coefficient (MCC) of 0.884, indicating a strong agreement between the automated and human filtering outcomes.
@@ 6.1 Experimental Setup
$\text{Call Acc}=\frac{\sum_{i=1}^{N}c_{i}}{\sum_{i=1}^{N}n_{i}}$ (1)
@@
To provide deeper insight, we generate a multi-turn counterpart of our dataset generated by standard self-instruct approach previously used in the literature (qin2023toolllm). We use the same goal APIs for generating short multi-turn tool-use instances. All necessary information to complete the task is revealed within a few turns, without any distractions.
@@ 6.2 Main Results
Closed-source models consistently outperform open-source models under noisy settings. While general-purpose open-source models (Qwen2.5, DeepSeek-V3) demonstrate more stable performance across scenarios, they still lag behind top proprietary models. Notably, open-source finetuned models like ToolACE and watt-tool achieve generally achieve poor accuracy of less than 10%, despite their high BFCL scores. We further examine this by extending the comparison between multi-turn and long-term tool-use performance in the following section.
@@ 6.4 Controlled Analysis on Long-Context Utilization of TALMs
Sample Dialogue from Missing Context-Hard:
(Previous sessions)
Ground-truth Response:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "MISSING"
GPT-4o-mini Prediction:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "3"
DeepSeek-V3 Output:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "1"
@@ 6.5 Error Analysis through the Lens of Hallucination
Closed and General-purpose models exhibit highly skewed error distributions, with over 75% of their failures attributable to Out-of-context hallucinations in MC-Easy and similarly high proportions in MC-Hard (63% and 77%). We find that several models tend to copy default values of missing parameter or filling out randomly generated strings without proper verification of contextual sufficiency. Finetuned TALMs, in contrast, show a different error distribution. In MC-Easy, TALMs reduce Out-of-context hallucinations to 43.39% and In-context hallucinations remain low (7.94%), while API errors account for nearly half (48.68%) of the total errors, suggesting that TALMs are more cautious about fabricating arguments outright from missing context (lower Out-of-context) but often misidentify the appropriate tool (API error), particularly under harder conditions. See Table 4 for detailed case studies illustrating model behaviors.
@@ 7 Conclusion
Acknowledgments
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (No.RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)), (No. RS-2024-00457882, National AI Research Lab Project), (2022-0-00077, RS-2022-II220077,AI Technology Development for Commonsense Extraction, Reasoning, and Inference from Heterogeneous Data). Jinyoung Yeo is the corresponding author.
@@ A.1 Tool-Augmented Language Models
Large Language Models (LLMs) augmented with external tools have emerged as a promising paradigm for overcoming inherent limitations in static parametric memory (komeili2022internet) and enabling real-time interaction with the external world (yang2024codewand; chae2025web). The term tool encompasses a broad range of functionalities, from simple calculators and Python interpreters (cai2024latm) to complex API and function calls (liu2024toolace). In this work, we focus specifically on the latter—LLMs interfacing with external APIs or invoking function calls to perform tasks.
@@
Early efforts concentrated on enabling LLMs to call tools via prompting and reasoning frameworks. For instance, parisi2022talm introduce the concept of Tool Augmented Language Models (TALM), demonstrating that an LLM can iteratively interact with non-differentiable tools through textual interfaces. Similarly, the ReAct framework (yao2022react) combined chain-of-thought reasoning with action directives, allowing an LLM to plan steps and invoke tools (e.g., web search or calculators) in a conversational loop. Subsequent research has shown that fine-tuning LLMs on tool-use data can significantly enhance their efficacy. Toolformer (schick2023toolformer), for example, is a fine-tuned model with self-generated tool API calls inserted into its training data, enabling it to decide when and how to use calculators, search engines, translators, and other APIs.
@@
With the rising demand for practical task-solving capabilities, recent proprietary LLMs have evolved to support tool or function calling as a core interface. Starting with GPT-4’s function calling API (openai_function_calling), OpenAI has progressively enhanced tool interoperability in their models, culminating in GPT-4.1 and GPT-4o, which natively support multimodal input, extended context windows, and robust tool invocation within conversation flows. Similarly, models like Claude 3.5 (anthropic_claude_3) and Gemini 1.5 (team2024gemini) offer built-in support for external API access and tool orchestration, enabling seamless integration of search, calculator, code execution, and retrieval systems in their agent pipelines. These developments mark a paradigm shift from pure text generation to real-world interaction, where tool use becomes a first-class citizen in the model’s interface. Alongside model-level enhancements, ecosystem-level tool agents have also proliferated. For example, OpenAI’s ChatGPT integrates plugin systems for third-party APIs and file browsing, while Google’s Bard/Gemini and Anthropic’s Claude provide tool-use scaffolding through memory and workflow APIs. These systems often combine planning, execution, and memory modules, highlighting the emerging need for robust, API-centric LLM benchmarks and evaluation methods.
@@
Recognizing the performance gap between open-source LLMs and proprietary models in tool calling, recent efforts have focused on equipping smaller or open LLMs with competitive tool-use abilities. For instance, Gorilla (patil2023gorilla) fine-tunes LLaMA models on a large corpus of API documentation and call traces, leveraging a retriever-generator architecture to improve accuracy and reduce hallucination. ToolLLM (qin2023toolllm) adopts a reasoning-guided planning strategy that explores multiple API invocation paths via a depth-first search over possible tool sequences. Salesforce’s xLAM (zhang2024xlam) provides a family of open LLMs specifically tuned for tool usage, achieving strong performance on multi-tool benchmarks. Other open initiatives such as ToolAlpaca (tang2023toolalpaca), AnyTool (anytool), and ToolBridge (jin2024toolbridge) introduce simulated training pipelines or meta-learning objectives for improving tool generalization in smaller models. These efforts aim to bridge the gap in real-world utility and provide a competitive open-source alternative to commercial tool-augmented LLM systems.